# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 6 Conclusions And Future Work
We show that an allocation is WEF-able only if its weighted envy graph does not contain positive cost cycles. Unlike the unweighted scenario ([15]), a given allocation that maximizes the utilitarian welfare across all reassignments of its bundles to agents is not a condition for WEF-able. The gap between the weighted setting and the unweighted setting in EF characterization raises the open question of other conditions under which weighted EF allocations may exist.
@@ A.1 Subsidy Bound of Algorithm 1
As Theorem 3.17 implies, Algorithm 1 computes a WEF-able allocation with a total subsidy of at most $(W-w_{1})V$. However, this bound is not tight. To understand why, consider the case of $2$ items, each valued at $V$ by agent $i\in\{1,\ldots,n-1\}$, who has an entitlement of $w_{i}\geq 2$, and $V-\epsilon$ by all other agents. Our algorithm will allocate all the items to agent $i$, resulting in a subsidy of $\frac{w_{j}}{w_{i}}2(V-\epsilon)$ by each other agent $j\neq i\in N$, leading to a total subsidy of $(W-w_{i})\frac{2(V-\epsilon)}{w_{i}}$, for arbitrarily small $\epsilon>0$.
@@ A.2 Subsidy Bound of Algorithm 3
As Theorem 5.8 implies, Algorithm 3 computes a WEF-able allocation with a total subsidy of at most $\frac{W}{w_{1}}-1$. However, this bound is not necessary tight.
First, there are cases where breaking ties in favor of the agent with the minimum entitlement leads to a WEF-able allocation with a better subsidy bound.
Example A.1.
Consider three agents with weights $w_{1}=1,w_{2}=2,w_{3}=3$ and three items. Agent 1 and agent 3 value all the items at 1, while agent 2 values all the items at 0.
the algorithm is executed as follows:
@@
    For $t=1$, since agent $2$ values all items at 0, the algorithm removes this agent from the game. Then, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1},\frac{1}{w_{3}}=\frac{1}{3}$. Consequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset,A_{2}^{1}=\emptyset$ and $A_{3}^{1}=\{o_{1}\}$.
@@
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$ and $\frac{v_{3}(A_{3}^{1})+1}{w_{3}}=\frac{2}{3}$. Subsequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{2}=\emptyset,A_{2}^{2}=\emptyset$ and $A_{3}^{2}=\{o_{1},o_{2}\}$.
  3. 3.
    For $t=3$, the algorithm compares $\frac{1}{w_{1}}=1$ and $\frac{v_{3}(A_{3}^{2})+1}{w_{3}}=\frac{3}{3}=1$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{3}=3$. As a result, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $3$, producing $A_{1}^{3}=\emptyset,A_{2}^{3}=\emptyset$ and $A_{3}^{3}=\{o_{1},o_{2},o_{3}\}$.
In the resulting allocation, agent 1 envies agent 3 by $1$, resulting in a subsidy of $1$ by agent 1. Agent 2 now envies agent 1 by 1 due to the subsidy, resulting in a subsidy of $2$ by agent 2. The total subsidy required is $3$. However, if at $t=3$ we allocate an item to agent 1, resulting in the allocation $A_{1}=\{o_{1}\},A_{2}=\emptyset,A_{3}=\{o_{2},o_{3}\}$, agent 3 envies agent 1 by $1-\frac{2}{3}=\frac{1}{3}$, which means the subsidy required by agent 3 is $1$. Agent 2 now envies agent 3 by $\frac{1}{3}$, resulting in a subsidy of $\frac{2}{3}$ by agent 2. The total subsidy required is $\frac{5}{3}<3$.
This is not yet clear in which cases it is preferable to break ties in favor of higher or lower entitlements.
Secondly, if agent 1, who has an entitlement of $w_{1}$, is the only agent who requires 0 subsidy, then the subsidy required by each other agent $j\neq 1\in N$ is not necessary bounded by $\frac{w_{j}}{w_{1}}$.
Example A.2.
Consider two agent with weight $w_{1}=2,w_{2}=3$ and two items, both values at 1 by both agents. the algorithm is executed as follows:
  1. 1.
    For $t=1$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1},\frac{1}{w_{3}}=\frac{1}{3}$. Consequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
  2. 2.
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{2}$ and $\frac{v_{2}(A_{2}^{1})+1}{w_{3}}=\frac{2}{3}$. Subsequently, the algorithm searches for the shortest transfer path starting at $2$ and ending at $i_{\text{0}}$, and finds the path $(2,i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $2$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{2}=\{o_{2}\}$ and $A_{2}^{2}=\{o_{1}\}$.
In the resulting allocation, agent 2 envies agent 1 by $\frac{1}{2}-\frac{1}{3}=\frac{1}{6}$, resulting in a subsidy of $\frac{1}{2}$ by agent 3. That is, agent 1 with lower entitlement is the only agent required 0 subsidy, but the subsidy required by agent 2 is smaller than $\frac{w_{2}}{w_{1}}=\frac{3}{2}$.
We still do not know how to tighten this bound further for the general case.

### v4

@@ 6 Conclusions And Future Work
We show that an allocation is WEF-able only if its weighted envy graph does not contain positive cost cycles. Unlike the unweighted scenario (halpern2019fair), a given allocation that maximizes the utilitarian welfare across all reassignments of its bundles to agents is not a condition for WEF-able. The gap between the weighted setting and the unweighted setting in EF characterization raises the open question of other conditions under which weighted EF allocations may exist.
@@ A.1 Subsidy Bound of Algorithm 1
As Algorithm 1 implies, Algorithm 1 computes a WEF-able allocation with a total subsidy of at most $(W-w_{1})V$. However, this bound is not tight. To understand why, consider the case of $2$ items, each valued at $V$ by agent $i\in\{1,\ldots,n-1\}$, who has an entitlement of $w_{i}\geq 2$, and $V-\epsilon$ by all other agents. Our algorithm will allocate all the items to agent $i$, resulting in a subsidy of $\frac{w_{j}}{w_{i}}2(V-\epsilon)$ by each other agent $j\neq i\in N$, leading to a total subsidy of $(W-w_{i})\frac{2(V-\epsilon)}{w_{i}}$, for arbitrarily small $\epsilon>0$.
@@ A.2 Subsidy Bound of Algorithm 3
As Section 5 implies, Algorithm 3 computes a WEF-able allocation with a total subsidy of at most $\frac{W}{w_{1}}-1$. However, with more careful analysis, we can prove a tighter bound.
There are two cases to consider:
@@
    Agent 1 with the minimum entitlement receives a positive subsidy. Together, Section 5 and Section 5 imply that $s_{i}\leq\frac{w_{i}}{w_{1}}$ for each agent $i\in N$. Since agent 1 does receive a positive subsidy, and by Section 3.1, there exists at least one agent who requires no subsidy, the total required subsidy is bounded by $\frac{W-w_{2}}{w_{1}}$.
@@
    Agent 1 with the minimum entitlement receives no subsidy. We can modify Section 5 in the following way: for each agent $i\notin R(t)$, where $t\in[T]$, $\ell_{i}(A^{t})\leq\frac{1}{w_{2}}$. By the proof of Section 5, $\ell_{i}(A^{t})\leq\ell_{i_{j+1}}(A^{t})$. If $i_{j+1}=i_{1}$, then $\ell_{i}(A^{t})\leq\ell_{i_{j+1}}(A^{t})\leq 0$ (because agent 1 requires no subsidy). Otherwise, $\ell_{i}(A^{t})\leq\ell_{i_{j+1}}(A^{t})\leq\frac{1}{w_{i_{j+1}}}\leq\frac{1}{w_{2}}$. Overall, the subsidy required by each agent is bounded by $\frac{w_{i}}{w_{2}}$, and by Section 3.1, there exists at least one agent who requires no subsidy. Therefore, the total required subsidy is bounded by $\frac{W-w_{1}}{w_{2}}$.
To sum up, the total required subsidy is at most $\max\Big{\{}\frac{W-w_{1}}{w_{2}},\frac{W-w_{2}}{w_{1}}\Big{\}}$.